# Estimating Body and Hand Motion in an Ego-sensed World

### v1

@@ A.3 Implementation details
We use a transformer [92] architecture with rotary positional embeddings [86] for EgoAllo’s denoising model $\mu_{\theta}(\vec{x}_{n},\vec{c},n)$. Latent representations $\vec{z}_{c}$ for conditioning sequences $\vec{c}$ are computed using six transformer blocks, each containing a self-attention layer followed by a 2-layer MLP. Skip connections are used throughout. The denoised output $\hat{\vec{x}}_{0}$ is produced by using six additional transformer blocks that take $\vec{x}_{n}$ as input, while conditioning on $\vec{z}_{c}$ via cross-attention. All hidden dimensions are set to 512. For guidance, we use infrastructure for sparse nonlinear least squares implemented in [100]. For additional details, we refer to our code release.

### v2

@@ A.3 Implementation details
We use a transformer [91] architecture with rotary positional embeddings [85] for EgoAllo’s denoising model $\mu_{\theta}(\vec{x}_{n},\vec{c},n)$. Latent representations $\vec{z}_{c}$ for conditioning sequences $\vec{c}$ are computed using six transformer blocks, each containing a self-attention layer followed by a 2-layer MLP. Skip connections are used throughout. The denoised output is produced by using six additional transformer blocks that take $\vec{x}_{n}$ as input, while conditioning on $\vec{z}_{c}$ via cross-attention. All hidden dimensions are set to 512. For guidance, we use infrastructure for sparse nonlinear least squares implemented in [100]. For additional details, we refer to our code release.
A.4 Runtimes
Without guidance, each denoising step takes around 0.05 seconds on an RTX 4090; our experiments use 30 DDIM steps for sampling. The guidance optimizer can typically converge in around 0.2 seconds on GPU. Runtimes are reported on length-128 sequences.